OpenStudy (anonymous):

sophmore algebra help? the volume of a cylinder V is given by the formula V=πr²h, where r is the radius and h is the height. a. solve the formula for h. b. find the height of a cylinder with volume 2500π feet and radius 10 feet.

OpenStudy (mathmale):

First of all, Culty, what does it mean to "solve the formula for h"?

OpenStudy (anonymous):

doesn't it mean find h?

OpenStudy (mathmale):

What have you done in the past in regard to solving a formula for one of the variables?

OpenStudy (anonymous):

sorry im reallllly bad at math

OpenStudy (anonymous):

i don't know. we took notes and stuff on this but i don't understand how to do this.

OpenStudy (mathmale):

In this context, what would "find h" mean? I'm not trying to give you a hard time; I would like to help you connect with work you've done before.

OpenStudy (anonymous):

i just thought it meant solve the equation to get h

OpenStudy (the_fizicx99):

I think its asking to solve for H or leave it on the right side alone

OpenStudy (the_fizicx99):

Hmm, I think its H = V/pi*r^2

OpenStudy (anonymous):

we did a problem together in class do you want me to write that one out and show you how we did it?

OpenStudy (the_fizicx99):

Sure

OpenStudy (anonymous):

no idea how you get that at all. just looks like a bunch of numbers

OpenStudy (the_fizicx99):

Ohhh, "Solve the formula for H" Meaning plug in the given values and solve for h

OpenStudy (anonymous):

the equation of a line containing the points (a,0) and (0,b) is given by the formula x/a + y/b = 1. a. solve the equation for y. (the answer is y=(1-x/a)b b. suppose the line contains the points (4,0), and (0,-2). if x = 3, find y. (answer is y= -1/2)

OpenStudy (anonymous):

that's the one we did together in class. i guess it's kinda like the one i'm asking

OpenStudy (the_fizicx99):

Hm, I'm not sure how you got that but try asking @SolomonZelman

OpenStudy (anonymous):

i just need help with the one i asked about

OpenStudy (mathmale):

You're given V=πr²h. You want to solve this equation for h. In other words, you want to isolate h. Write V=πr²h and divide both sides by πr². The result is your formula for h. That's it.

OpenStudy (the_fizicx99):

O.O That what I did but it said "solve for h" I thought it meant 2 things, isolate h or solve for the height with the given values lol

OpenStudy (anonymous):